This block is at position (1,3), but because the optional argument [0.5,0.5] has been given, it is the centre of the block which is located at that point, rather than the topleft corner.

Here is another, slightly narrower, block, at position (2,1) on the page.

This block is 3 modules wide, and is placed with its top left corner at the ‘origin’ on the page. Note that the length of the block is not specified in the arguments – the box will be as long as necessary to accomodate the text inside it. You need to examine the output of the text to adjust the positioning of the blocks on the page.